A store wants to ensure a shelf full of marshmallow peeps as the holiday season approaches. Daily demand for peeps is normally distributed with a mean of 25 and a standard deviation of 5. Lead time is 3 days and the store intends a 90% service level.

What is their reorder point?
Answer #1

Reorder point in case of variable demand:

R = d * LT + z(Sigma(standard deviation) * LT)

Z value for a 10% stockout change is 1.282


R = 25 * 3 + (1.282 * 5 * 3)

Reorder point = 75 + 19.23

ROP = 94.23 = 94